OpenStudy (anonymous):

Ken has a pole that is 26 inches long. He wishes to cut it into two pieces so that one piece will be 8 inches longer than the other. How long must the shorter piece be?

OpenStudy (anonymous):

Hi tendernilla, I can help you out with this one! We will have to set up an equation to solve this question.

OpenStudy (anonymous):

We are two different pieces, one of which has a length defined by the other piece. Using the variable x, how do you think we should express these lengths?

OpenStudy (blackbird02):

let x be the shorter piece x+8 be the longer piece

OpenStudy (anonymous):

I think you should let the original poster try their hand at answering the question

OpenStudy (blackbird02):

The equation would be x+(x+8)=26 So, x+(x+8)=26 2x+8 = 26 2x = 26 - 8 2x = 18 x = 18/2 x = 9

OpenStudy (blackbird02):

The shorter piece is 9 inches long.